OpenStudy (aryana_maria2323):

Please help. Will FAN AND MEDAL! Find the average value of f(x)=e2x over the interval [2, 4]. a. 1463.18 b. 731.59 c. 1517.78 d. 23.60

satellite73 (satellite73):

average value is is \[\frac{1}{b-a}\int_a^bf(x)dx\]

satellite73 (satellite73):

in your case \[\frac{1}{2}\int_2^4e^xdx\]

OpenStudy (aryana_maria2323):

Okay so I just plug that into my calculator and I will give me the answer? @Satellite73

satellite73 (satellite73):

just integrate it it will be \[\frac{1}{2}\left(e^4-e^2\right)\]

OpenStudy (aryana_maria2323):

Okay so I replace e^4 with 2^4 and e^2 with 4^2 right?

satellite73 (satellite73):

i am sorry i told you wrong

satellite73 (satellite73):

you have to find the anti derivative first

satellite73 (satellite73):

\[\frac{1}{2}\int_2^4e^{2x}dx\]

satellite73 (satellite73):

i put \(e^x\) that was a mistake

satellite73 (satellite73):

you know the anti derivative of \(e^{2x}\)?

OpenStudy (aryana_maria2323):

\[f(x)= \frac{ e^2x }{ 2 }+c\]

OpenStudy (aryana_maria2323):

@Satellite73

OpenStudy (aryana_maria2323):

@sleepyjess

OpenStudy (aryana_maria2323):

Can you please help? @sleepyjess told me that you might be able to @perl @TheSmartOne

OpenStudy (aryana_maria2323):

Do you have any ideas? @MrNood

OpenStudy (mrnood):

the formula you wrote last is the correct integral so work out the definite integral between 2 and 4 then divide by 4-2

OpenStudy (aryana_maria2323):

Which formula? The anti-derivative one?

OpenStudy (mrnood):

the formula you wrote last

OpenStudy (mrnood):

substitute 4 into your equation substitute 2 into your equation subtract those 2 numbers divide by 4-2

OpenStudy (aryana_maria2323):

In the question it has [2,4]. Am I subtracting 2-4 or 4-2 and then dividing by 2-4 or 4-2? I only ask because in the problem the 2 comes before the 4. @MrNood

OpenStudy (perl):

a = 2, b = 4 b-a = 4-2

OpenStudy (aryana_maria2323):

Okay so in the equation and I replacing the e or the x with 2 and 4?

OpenStudy (aryana_maria2323):

The x right?

OpenStudy (perl):

\[ f_{avg} = \frac {1}{b-a} \int_{a}^{b} f(x) dx \]\[ f_{avg} = \frac {1}{4-2} \int_{2}^{4} e^{2x} dx = \frac 1 2 \int_{2}^{4} e^{2x} dx \] Now you just need to evaluate the integral and multiply by 1/2.

OpenStudy (aryana_maria2323):

What I don't understand?

OpenStudy (mrnood):

you haven't made any attempt to actually DO the process substitute for X and do what I said in my last post

OpenStudy (aryana_maria2323):

I have attempted. I substituted for x and got this. \[f(2)=\frac{ e^2x }{ 2 }+c = e^2+c\] \[f(4)=\frac{ e^4x }{ 2 }+c = 2e^2+c\] \[\frac{2e^2+c-e^2+c }{ 4-2 }= \frac{ 1 }{ 2 }\] I am not 100% sure I am correct but this is what I understood from what you said. @MrNood @perl

OpenStudy (mrnood):

you have misread the equation it is e^(2x) not e^2 (x) you do not need the constant - since this is a definite integral \[\frac{ e ^{2x} }{2 }\] substitute x=4 substitute x= 2 subtract those 2 numbers divide by 4-2

OpenStudy (aryana_maria2323):

am I finding the derivative of when I substitute for x or am I just evaluating/simplifying the problem?

OpenStudy (aryana_maria2323):

After doing that I got B. 731.59?

OpenStudy (aryana_maria2323):

Thank you @MrNood